OpenStudy (anonymous):

The table below shows the values of y for each value of x: x y -1 0 -1 -1 -2 2 -2 3

OpenStudy (anonymous):

 lol sorry chart sucks

OpenStudy (imstuck):

what are you doing with the values?

OpenStudy (anonymous):

Part A: Does the table represent a relation that is a function? Justify your answer by using the values shown in the table. (4 points) Part B: The function f(h) shown below represents the number of tennis matches Diana won each month when she practiced tennis for h hours every day: If f(h) = 3h + 1 Calculate and interpret the meaning of f(3). (4 points) Part C: Write an ordered pair to represent the input and output of the function in Part B when Diana practices tennis for b hours every day. (2 points)

OpenStudy (anonymous):

sorry i know its a lot

OpenStudy (solomonzelman):

\(\huge\color{ blue }{\huge {\bbox[5pt, cyan ,border:2px solid white ]{ \LARGE{ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ }\\ \begin{array}{|c|c|c|c|} \hline~~\textbf{x}~~~~~~ ~&~~~~~~\textbf{f(x) }~~~~~~\\ \hline\text{-1} &\text{0} \\ \text{-1} &\text{-1} \\ \text{-2} &\text{2} \\ \text{-2} &\text{3} \\ \hline \end{array} }}}\)

OpenStudy (solomonzelman):

with -1 you get 0 and -1. This is only possible in a case of `x=-1` But the rest x's tell em that there are Xs that are equal to -2. No it is not a function.

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(h) = 3h + 1}\) \(\normalsize\color{blue}{ f(3) = 3(3)+ 1~~~~→~~~f(3)=7}\)

OpenStudy (anonymous):

it keeps saying math processing error

OpenStudy (solomonzelman):

f(3) gives the value, for the matches won, per 3 hours. So she won 7 matches when she practiced 3 hours per day.

OpenStudy (solomonzelman):

what keeps saying math processing error ?

OpenStudy (anonymous):

idk it looks like you were saying something and then it said math processing error

OpenStudy (solomonzelman):

f(h) = 3h + 1 for part C (playing b hours every day, it should just be f(b) = 3b + 1 .

OpenStudy (solomonzelman):

CAn you see everything clearly, or do you see some not-making sense codes like `\(\normalsize\color{blue}{ (h) = 3h + 1}\) `

OpenStudy (anonymous):

for the above?

OpenStudy (solomonzelman):

I mean when I posted colored functions/ numbers ...

OpenStudy (solomonzelman):

you see everything just fine ?

OpenStudy (anonymous):

its not a code or anything it just says math processing error

OpenStudy (anonymous):

and what would the ordered pair be for part C?

OpenStudy (solomonzelman):

well, we don't know the b, so it would just be f(b)=3b+1

OpenStudy (anonymous):

i think its a typo and they ment to put h cuz they both mean hours per day

OpenStudy (solomonzelman):

well, we don't knwo the h either, do we? We can say f(h)=3h+1 (to just copy our initial function)

OpenStudy (solomonzelman):

the domain would be [0,∞) the range would be [1,∞)

OpenStudy (anonymous):

i thought the h was 3?

OpenStudy (solomonzelman):

Wasn't that before C?